# Do machines fail like humans? A human-centred out-of-distribution spectrum for mapping error alignment

### 08 Mar 2026 v1

@@ Abstract
Determining whether AI systems process information similarly to humans is central to cognitive science and trustworthy AI. While modern AI models match human accuracy on standard tasks, such parity does not guarantee that their underlying decision-making strategies are aligned with human information processing. Assessing performance using i) error alignment metrics to compare how humans and models fail, and ii) using distorted, or otherwise more challenging, stimuli, provides a viable pathway toward a finer characterization of model-human alignment. However, existing out-of-distribution (OOD) analyses for challenging stimuli are limited due to methodological choices: they define OOD shift relative to model training data or use arbitrary distortion-specific parameters with little correspondence to human perception, hindering principled comparisons. We propose a human-centred framework that redefines the degree of OOD as a spectrum of human perceptual difficulty. By quantifying how much a collection of stimuli deviates from an undistorted reference set based on human accuracy, we construct an OOD spectrum and identify four distinct regimes of perceptual challenge. This approach enables principled model-human comparisons at calibrated difficulty levels. We apply this framework to object recognition and reveal unique, regime-dependent model-human alignment rankings and profiles across deep learning architectures. Vision-language models are the most consistently human aligned across near- and far-OOD conditions, but CNNs are more aligned than ViTs for near-OOD and ViTs are more aligned than CNNs for far-OOD conditions. Our work demonstrates the critical importance of accounting for cross-condition differences such as perceptual difficulty for a principled assessment of model-human alignment.
@@
Keywords Error alignment $\cdot$ Human-centred OOD
@@ 1 Introduction
Evaluating whether artificial intelligence systems process information in ways comparable to humans is a fundamental challenge for both cognitive science and computer science. From a cognitive science perspective, neural network models offer computational abstractions of hypothesised mechanisms underlying human information processing [23, 4, 32], and assessing their alignment with humans provides a systematic way to test these hypotheses [34, 36]. From a computer science perspective, models whose information processing reflects human cognition may exhibit greater robustness [31, 30], interpretability [10, 3], and trustworthiness [26, 19]. As humans generalise remarkably well across diverse and imperfect environments, models that better reflect human processing principles may inherit similar robustness properties. Existing AI models achieve or exceed human accuracy on clear inputs. However, human-level accuracy does not entail model-human alignment, as systems may make markedly different errors reflecting different processing mechanisms. Thus, going beyond accuracy is crucial [2, 34]. Metrics such as error consistency (EC)[11] and misclassification agreement (MA)[44] address this by measuring error patterns between systems. Yet comparing errors on easy tasks is insufficient – alignment may appear high on clear stimuli simply due to the infrequency of errors. Challenging conditions, where error patterns can become more distinct, are essential for rigorously assessing model-human alignment.
@@
Although these considerations apply broadly, visual perception provides a particularly well-defined testbed for studying alignment under challenging conditions. When systems are exposed to degraded visual input, errors reveal inductive biases – the visual features they degrade reflecting their information-processing strategies  [13, 1, 5]. Distortions that selectively disrupt edges, texture, global structure, or high-frequency detail differentially impair systems depending on which cues they privilege (e.g., texture or global shape). Researchers have sought to compare the behavioral patterns of humans and models under stimuli of different distortion types, which they call ‘out-of-distribution’ (OOD) samples  [22, 7, 14, 17]. However, to assess model-human alignment appropriately, performance must be compared on a common and meaningful scale. We identify four fundamental problems for model-human comparison under distorted conditions and provide a human-centered behavioral deviation framework to solve these issues.
Figure 1: Example images from two distortion types (low-pass and high-pass filtering) from the modelvshuman datase [12] at different parameter levels. Images within the same column share the same parameter rank within each distortion type.
@@
First, OOD has been defined based on models rather than humans. In machine learning, OOD is defined by deviation from training data statistics [9]. Recent work focuses on OOD samples for models [18, 17, 39]. However, for humans, this definition does not straightforwardly apply. Humans do not have a finite, controlled "training distribution," but instead acquire visual knowledge through a lifetime of varied, naturally encountered experience. This asymmetry raises fundamental questions: Are OOD samples for models also OOD for humans? For instance, is Figure 1, OOD for humans? What does OOD mean for human perception, and how should it be defined?
@@
Second, the level of degradation is typically defined by image-processing pipeline parameters rather than a universal measure. Consider the distortions in Figure 1. If we want to compare the impact of specific feature impairment, which conditions should we choose? Should we select those within the same column? If human perceptual difficulty varies across conditions it is unclear if differences in error patterns reflect feature or information-processing differences or difficulty. A low-pass filter with parameter 5 and a high-pass filter with parameter 1 cannot be meaningfully placed on a common scale, despite sharing the same parameter rank. Some studies aggregate performance across severity levels  [17, 12, 44], but such an approach ignores how perceptual difficulty varies across a distortion type. Others qualitatively select distortion levels that appear roughly comparable  [6, 7, 43, 29], but this introduces subjective judgement. Hendrycks et al.  [17] normalised to a reference model (AlexNet [24]), but this substitutes one arbitrary reference for another where the resulting scale reflects the idiosyncrasies of that architecture rather than any principled notion about image degradation.
@@
Third, not all degradation conditions are equally meaningful for the investigation. Consider an extreme example where an image with a low-pass parameter of 40 appears unrecognisable to humans (Figure 1) – would such a stimulus be appropriate for assessing model-human alignment? If not, how can we determine the appropriate stimulus set for each distortion type?
@@
Lastly, raw model-human alignment values can be misleading without a baseline. Previous work treats model-human alignment values directly as evidence of alignment [12]. However, this interpretation is incomplete without considering a baseline such human–human alignment values. When humans show zero agreement on a stimulus set, we cannot expect high model-human alignment.
@@
To address these challenges, we propose a human-centred behavioral deviation framework that provides a natural benchmark for fairly comparing model-human alignment across distortion types on a common scale. Rather than using distortion-specific parameters (Figure 1), we place them on a shared scale anchored in human perceptual difficulty (Figure 2B). We operationalise this using the effect size Glass’s $\Delta$ [16] applied to accuracy logits – the deviation of human performance from undistorted stimuli, which we term the reference distribution. We refer to this quantity as the OOD score. With this, we redefine OOD for humans as deviation from this reference condition that impairs human perception in quantifiable, graded ways. Moreover, for cognitive science, defining OOD in terms of human perceptual difficulty provides a principled way to control task difficulty across conditions, enabling systematic investigation of individual differences in visual processing strategies that would be confounded by arbitrary distortion parameters [21].
@@
Our analysis is conducted on the modelvshuman dataset [12], which contains human core object recognition performance on systematically distorted images. To preview the results, our approach reveals that a perceptual OOD spectrum contains distinct regions that reflect different regimes of human information processing. We evaluate models within near- and far-OOD regions, showing distinct model-human alignment patterns across regimes: Vision-Language Models (VLMs) [33] show the most consistent alignment across regimes, whereas Vision Transformer models (ViTs) [8] achieve higher alignment than Convolutional Neural Networks (CNNs) [35] in near-OOD but the reverse pattern for far-OOD. In sum, this work introduces a human-centred framework to appropriately compare human and model error profiles, and offer an empirical characterisation of how different model architectures align with human perception across regimes.
@@ 2 Results
Figure 2: A human-centered OOD spectrum. (A) Illustration of the process for measuring the degree of human performance deviation on a specific distortion condition relative to undistorted images. (B) OOD spectrum constructed from human accuracy (logit transformed). The x-axis represents the OOD score, indicating the degree of deviation from the performance baseline on undistorted images, with greater negative values corresponding to larger distributional shifts. Example images from selected distortions are shown above their corresponding positions on the axis. (C) OOD scores for all levels of corruption for each distortion type in modelvshuman dataset. (D) Four-component Gaussian Mixture Model (GMM) fitted to OOD scores across all distortion conditions, defining the four OOD groups. Color reflects different OOD regimes. Boundaries illustrated by dashed lines.
@@ 2.1.2 Building the OOD Spectrum through Quantifying Human Performance Deviation
To build the perceptual OOD spectrum, we construct a reference distribution that quantifies how much each distorted condition deviates from the undistorted baseline (i.e., the effect size), providing a continuous measure of distortion severity grounded in human performance. Due to non-normality in the accuracy scores (Shapiro–Wilk test, $p<0.05$), we applied a logit transformation to the scores, after which all tests yielded $p>0.05$, indicating no rejection of the null hypothesis of normality (see Appendix A, Table S4).
@@
where $\bar{l_{d}}$ and $\bar{l_{ud}}$ are the mean logit-transformed accuracies on the distorted set and undistorted set respectively, and $s_{ud}$ is the standard deviation for the logit-transformed accuracy on the undistorted images. Glass’s $\Delta$ is appropriate here because it standardises the mean difference between performance on distorted and undistorted images using only the reference condition’s variance, ignoring the variance of each distorted condition. This ensures comparability across conditions by measuring all deviations on a common scale defined by the reference – the baseline condition – and avoids underestimating behavioral differences for distortion conditions with high variance.
@@ 2.1.3 The Perceptual Distortion OOD Spectrum
To meaningfully assess model robustness in relation to human perception, we construct a human-centred spectrum OOD based on behavioral deviation across distortion types and levels, which allows us to replace the arbitrary distortion parameter scales. To illustrate the range of behavioral deviation, we present three example distortion types (Figure 2.B). The behavioral deviation effect sizes for all distortion conditions are presented in Figure 2.C. To explore error profiles and investigate the distinct inductive biases by different distortion types, we used a data-driven approach to group conditions that produce comparable degrees of behavioral deviation. We fitted a Gaussian Mixture Model (GMM) to the distribution of OOD scores across all distortion types and levels (including the undistorted sets), which consistently gave a four-component solution (Bayesian Information Criterion (BIC) and Akaike Information Criterion (AICc); Figure S8 in Appendix. A).
@@ 2.2.1 Error alignment across OOD regimes are markedly different
We assess human-human error alignment across the four OOD regimes (Figure 3A). We find that extreme-OOD distortions consistently yielded much lower EC values compared to near- and far-OOD conditions, highlighting a pronounced breakdown in shared perceptual strategies under severe distortion (Figure 3A, top). As noted, all conditions in the extreme-OOD region are consistent with chance level accuracies. Distortions also differed in the rate at which EC declines across OOD levels. For instance, uniform noise and phase scrambling show a sharp drop of EC in the far-OOD regime, whereas distortions such as high-pass decrease more gradually, likely reflecting differences in how these manipulations disrupt perceptual features.
Similar to EC, MA declines with increasing perceptual difficulty, with extreme-OOD levels consistently producing the lowest scores (Figure 3A, bottom). In contrast, the difference between far-OOD and extreme-OOD is less pronounced than EC for some distortion types such as eidolon-I, II and high-pass. The ordering of alignment values across OOD regimes from near to extreme mirrors EC closely, reinforcing the robustness of our perceptual difficulty–based spectrum. The violin plots complement these findings by showing that MA distributions are generally more concentrated than EC distributions and are typically centred at lower values. This indicates that, while participants may often make errors on the same items (captured by EC), it is far less common for them to predict the same incorrect class.
2.2.2 Human error profiles are primarily structured by perceptual difficulty (OOD level) rather than distortion type
@@
Previous studies have often treated each distortion type (e.g., uniform noise or low-pass filtering) as a monolithic condition. Nonetheless, distortions that differ in visual form can impose comparable perceptual challenges and elicit similar error profiles. To capture these relationships, we define each domain as a specific distortion type at a given severity level and construct a CLED-based distance matrix across all domains. This analysis allows us to assess the relative contributions of distortion type and OOD severity to the structure of human error behavior.
The similarity of error profiles across human participants is visualized using t-SNE (Figure 3B). If different severity levels within a distortion type produced similar error profiles, we would expect tight, coherent clusters for each distortion type – conditions of the same type would group together regardless of severity. However, this is not the case: lower and higher severity levels within the same distortion type can be far apart, while conditions from different distortion types sometimes appear more closely aligned. Notably, conditions within the near-OOD regime (green labels) cluster tightly together, whereas their counterparts in far- (blue) or extreme-(purple) OOD regimes are typically more dispersed.
We confirm this through quantitative analyses comparing the effect of distortion type and level. We estimate the separability of within- versus between-group CLED values for distortion-type-based grouping and OOD-level-based grouping using Cohen’s $D$ and performed permutation tests for statistical inference. The negative values indicate that error patterns of samples within a group (e.g., an OOD level or a distortion type) are more internally consistent than between groups (Table 1). Across both measures, OOD level consistently exhibited a much stronger effect than distortion type, reinforcing that perceptual difficulty is a crucial factor structuring human error profiles.
@@
Human behavior provides a foundation for model–human comparison: the degree of human-human alignment establishes an empirical upper bound on the alignment a model can achieve. Here, we examine model–human alignment across OOD regimes and which architectures show more human-like error patterns.

### 11 Apr 2026 v2

@@ Abstract
Determining whether AI systems process information similarly to humans is central to cognitive science and trustworthy AI. While modern AI models can match human accuracy on standard tasks, such parity does not guarantee that their underlying decision-making strategies resemble those of humans. Assessing performance using error alignment metrics to compare how humans and models fail, and how this changes for distorted, or otherwise more challenging, stimuli, provides a viable pathway toward a finer characterization of model-human alignment. However, existing out-of-distribution (OOD) analyses for challenging stimuli are limited due to methodological choices: they define OOD shift relative to model training data or use arbitrary distortion-specific parameters with little correspondence to human perception, hindering principled comparisons. We propose a human-centred framework that redefines the degree of OOD as a spectrum of human perceptual difficulty. By quantifying how much a collection of stimuli deviates from an undistorted reference set based on human accuracy, we construct an OOD spectrum and identify four distinct regimes of perceptual challenge. This approach enables principled model-human comparisons at calibrated difficulty levels. We apply this framework to object recognition and reveal unique, regime-dependent model-human alignment rankings and profiles across deep learning architectures. Vision-language models are most consistently human aligned across near- and far-OOD conditions, but convolutional neural networks (CNNs) are more aligned than vision transformers (ViTs) for near-OOD and ViTs are more aligned than CNNs for far-OOD. Our work demonstrates the critical importance of accounting for cross-condition differences, such as perceptual difficulty, for a principled assessment of model-human alignment.
@@
Keywords Model–human alignment $\cdot$ Error alignment $\cdot$ Out-of-distribution (OOD) robustness $\cdot$ Computer vision
@@ 1 Introduction
Evaluating whether AI systems process information like humans is a core challenge for cognitive and computer science. Neural networks can serve as computational abstractions to test hypotheses of human processing in cognitive science [23, 4, 32, 34, 36], while in computer science, such alignment improves robustness [31, 30], interpretability [10, 3], and trustworthiness [26, 19]. Existing AI models achieve or exceed human accuracy on clear inputs. However, human-level accuracy does not entail model-human alignment, as systems may make different errors reflecting completely different processing mechanisms. Thus, going beyond accuracy is crucial [2, 34]. Metrics such as error consistency (EC)[11] and misclassification agreement (MA)[44] address this by measuring error patterns between systems. Yet comparing errors on easy tasks is insufficient – alignment may appear high on clear stimuli simply due to the infrequency of errors. Challenging conditions, where error patterns can become more distinct, are essential for rigorously assessing model-human alignment.
@@
In previous alignment analyses of visual perception, degraded inputs have revealed inductive biases and information-processing strategies [13, 1, 5]. Research comparing human and model behavior under ’out-of-distribution’ (OOD) samples [22, 7, 14, 17] suggests that distortions to features like texture or global shape differentially impair systems based on the cues they privilege. However, to assess model-human alignment appropriately, performance must be compared on a common and meaningful scale. We identify four fundamental problems for model-human comparison under distorted conditions and provide a human-centred behavioral deviation framework to solve these issues.
@@
First, there is no clear meaning to OOD stimuli in humans. In machine learning, OOD is defined by deviation from training data statistics [9]. Recent work on model-human alignment is guided by this definition [18, 17, 39], but humans do not have a finite, controlled “training distribution”, and instead acquire knowledge through a lifetime of varied, naturally encountered experience. This asymmetry raises fundamental questions: How should we define OOD for human perception? For instance, which images in Figure 1A are OOD for humans?
Figure 1: A human-centered OOD spectrum. (A) Example images from two distortion types (left column: low-pass filtering; right column: high-pass filtering) from the modelvshuman datase [12] at different parameter levels. Images within the same row share the same parameter rank within each distortion type. (B) Illustration of the process for measuring the degree of human performance deviation on a specific distortion condition relative to undistorted images. (C) OOD scores for all levels of corruption for each distortion type in modelvshuman dataset. The x-axis represents the OOD score, indicating the degree of deviation from the performance baseline on undistorted images, with greater negative values corresponding to larger distributional shifts. (D) Four-component Gaussian Mixture Model (GMM) fitted to OOD scores across all distortion conditions, defining the four OOD groups. Color reflects different OOD regimes. Boundaries illustrated by dashed lines. (E) OOD spectrum with example images from selected distortions.
@@
Second, there is no unified principled measure of degree of distortion in OOD stimuli. Consider the distortions in Figure 1A, how do we compare the degree of distortion between images in the first row and those in the second? When investigating model-human alignment in distorted images, some studies aggregate performance across severity levels  [17, 12, 44], but such an approach ignores how perceptual difficulty varies across a distortion type. Others qualitatively select distortion levels that appear roughly comparable  [6, 7, 43, 29], but this introduces subjective judgement. Hendrycks et al.  [17] identified the problem with grouping distortion conditions with different difficulties and address this by normalising to a reference model, but this substitutes one arbitrary reference for another where the resulting scale reflects the idiosyncrasies of that architecture rather than any principled notion about image degradation.
@@
Third, not all degradation conditions are equally meaningful (as raised in [12]). Consider an extreme example where an image with a low-pass parameter of 40 appears unrecognizable to humans (Figure 1A) – would such a stimulus be appropriate for assessing model-human alignment? If not, how should we determine the appropriate stimulus set?
@@
Lastly, raw model-human alignment values lack meaning without a baseline. Previous work treats alignment values directly as evidence of alignment [12]. However, this interpretation is incomplete without considering a baseline of such human–human alignment values. When humans show zero agreement on a stimulus set, we cannot expect high model-human alignment.
@@
To address these challenges, we propose a human-centred behavioral deviation framework that provides a natural benchmark for fairly comparing model-human alignment across distortion types on a common scale. Rather than using distortion-specific parameters (Figure 1A), we place them on a shared scale anchored in human perceptual difficulty (Figure 1C). We operationalise this OOD score using the effect size Glass’s $\Delta$ [16] applied to accuracy logits – deviation of human performance from undistorted stimuli, which we term the reference distribution. With this, we define OOD for humans as deviation from the reference condition, providing a quantitative scale we call the perceptual OOD spectrum. Defining OOD in terms of human perceptual difficulty provides a principled way to control task difficulty across conditions, enabling systematic investigation of visual processing strategies that would be confounded by arbitrary distortion parameters [21].
@@
Our analysis is conducted on the modelvshuman dataset [12], which contains human core object recognition performance on systematically distorted images. To preview the results, our approach reveals that a perceptual OOD spectrum contains distinct regions that reflect different regimes of human information processing. We evaluate models within near- and far-OOD regions, showing distinct model-human alignment patterns across regimes: Vision-Language Models (VLMs) [33] show the most consistent alignment across regimes, whereas Vision Transformer models (ViTs) [8] achieve higher alignment than Convolutional Neural Networks (CNNs) [35] in near-OOD but the reverse pattern for far-OOD. However, despite strong alignment in some models, human–human alignment remains a ceiling that no model approaches, indicating that the visual features humans rely on overlap more with each other than with those exploited by models. In sum, this work introduces a human-centred framework to appropriately compare human and model error profiles, and offer an empirical characterization of how different architectures align with human perception across regimes.
@@ 2.1.2 Building the OOD Spectrum through Quantifying Human Performance Deviation
To build the perceptual OOD spectrum, we construct a reference distribution that quantifies how much each distorted condition deviates from the undistorted baseline (i.e., the effect size), providing a continuous measure of distortion severity grounded in human performance. Due to non-normality in the accuracy scores (Shapiro–Wilk test, $p<0.05$), we applied a logit transformation, after which all tests yielded $p>0.05$, indicating no rejection of the null hypothesis of normality (see Appendix A, Table S4).
@@
where $\bar{l_{d}}$ and $\bar{l_{ud}}$ are the mean logit-transformed accuracies on the distorted set and undistorted set respectively, and $s_{ud}$ is the standard deviation for the logit-transformed accuracy on the undistorted images. Glass’s $\Delta$ is appropriate because it standardises the mean difference between performance on distorted and undistorted images using only the reference condition’s variance, ignoring the variance of each distorted condition. This ensures comparability across conditions by measuring all deviations on a common scale defined by the reference – the baseline condition – and avoids underestimating behavioral differences for distortion conditions with high variance.
@@ 2.1.3 The Perceptual Distortion OOD Spectrum
To meaningfully assess model robustness in relation to human perception, we construct a human-centred OOD spectrum based on behavioral deviation across distortion types and levels, to replace the arbitrary distortion parameter scales. To illustrate the range of behavioral deviation, we present three example distortion types (Figure 1E). The behavioral deviation effect sizes for all distortion conditions are presented in Figure 1C. To explore error profiles and investigate the distinct inductive biases by different distortion types, we used a data-driven approach to group conditions that produce comparable degrees of behavioral deviation. We fitted a Gaussian Mixture Model (GMM) to the distribution of OOD scores across all distortion types and levels (including the undistorted sets), which consistently gave a four-component solution (using Bayesian Information Criterion (BIC) or Akaike Information Criterion (AICc); see Figure S 7).
We interpret these groups as qualitatively different regimes of perceptual difficulty (Figure 1D). We label these groups as reference, near-OOD, far-OOD, and extreme-OOD, ordered by group means. Notably, the conditions that fall within the extreme-OOD level correspond precisely to those yielding chance level human performance, suggesting little or no recognisable information and are potentially of less interest than the other OOD regimes.
@@ 2.2.1 Error alignment across OOD regimes are markedly different
We assess human-human error alignment across the four OOD regimes (Figure 2A). We find that extreme-OOD distortions consistently yielded much lower EC values compared to near- and far-OOD conditions, highlighting a pronounced breakdown in shared perceptual strategies under severe distortion (Figure 2A, top). As noted, all conditions in the extreme-OOD region are consistent with chance level accuracies. Distortions also differed in the rate at which EC declines across OOD levels. For instance, uniform noise and phase scrambling show a sharp drop of EC in the far-OOD regime, whereas distortions such as high-pass decrease more gradually. Similar to EC, MA declines with increasing perceptual difficulty, with extreme-OOD levels consistently producing the lowest scores (Figure 2A, bottom). In contrast, the difference between far-OOD and extreme-OOD for MA is less pronounced than for EC in some distortion types such as eidolon and high-pass. The ordering of alignment values across OOD regimes from near to extreme mirrors EC closely, reinforcing the robustness of our perceptual difficulty–based spectrum. The violin plots complement these findings by showing that MA distributions are generally more concentrated than EC distributions and are typically centred at lower values.
2.2.2 Human error profiles are primarily structured by perceptual difficulty (OOD score)
@@
Previous studies have often treated each distortion type (e.g., uniform noise or low-pass filtering) as a monolithic condition [14]. Nonetheless, distortions that differ in visual form can impose comparable perceptual challenges and elicit similar error profiles. To capture these relationships, we define each domain as a specific distortion type at a given severity level and construct a CLED-based distance matrix across all domains. This analysis allows us to assess the relative contributions of distortion type and OOD severity to the structure of human error behavior.
The similarity of error profiles across human participants is visualized using t-SNE (Figure 2B). If different severity levels within a distortion type produced similar error profiles, we would expect tight, coherent clusters for each distortion type – conditions of the same type would group together regardless of severity. However, this is not always the case: lower and higher severity levels within the same distortion type can be far apart, while conditions from different distortion types sometimes appear more closely aligned. Notably, conditions within the near-OOD regime (green labels) cluster tightly together, whereas their counterparts in far- (blue) or extreme-(purple) OOD regimes are typically more dispersed.
We confirm this through quantitative analyses comparing the effect of distortion type and level. We estimate the separability of within- versus between-group CLED values for distortion-type-based grouping and OOD-level-based grouping using Cohen’s $D$ and performed permutation tests for statistical inference. The negative values indicate that error patterns of samples within a group (e.g., an OOD level or a distortion type) are more internally consistent than between groups (Table 1). Across both measures, OOD level consistently exhibited a much stronger effect than distortion type, reinforcing that perceptual difficulty is a crucial factor influencing human error profiles.
@@
Human behavior provides a foundation for model–human comparison: the degree of human-human alignment establishes an empirical upper bound on achievable model-human alignment. Here, we examine model–human alignment across OOD regimes and which architectures show more human-like error patterns.